# Institutional AI: A Governance Framework for Distributional AGI Safety

### v1

@@ Abstract
As LLM-based systems increasingly operate as agents embedded within human social and technical systems, alignment can no longer be treated as a property of an isolated model, but must be understood in relation to the environments in which these agents act. Even the most sophisticated methods of alignment, such as Reinforcement Learning through Human Feedback (RHLF) or through AI Feedback (RLAIF) cannot ensure control once internal goal structures diverge from developer intent. We identify three structural problems that emerge from core properties of AI models: (1) behavioral goal-independence, where models develop internal objectives and misgeneralize goals; (2) instrumental override of natural-language constraints, where models regard safety principles as non-binding while pursuing latent objectives, leveraging deception and manipulation; and (3) agentic alignment drift, where individually aligned agents converge to collusive equilibria through interaction dynamics invisible to single-agent audits. The solution this paper advances is Institutional AI: a system-level approach that treats alignment as a question of effective governance of AI agent collectives. We argue for a governance-graph that details how to constrain agents via runtime monitoring, incentive shaping through prizes and sanctions, explicit norms and enforcement roles. This institutional turn reframes safety from software engineering to a mechanism design problem, where the primary goal of alignment is shifting the payoff landscape of AI agent collectives.
@@
This paper offers the first full-length exposition of the Institutional AI framework: a theoretical approach to alignment that relocates safety guarantees from training-time internalization to runtime institutional structures. A companion paper, Institutional AI: Governing LLM Collusion in Multi-Agent Cournot Markets via Public Governance Graphs, applies the framework to a concrete problem: preventing autonomous pricing agents from learning to collude against consumer welfare in a multi-agent Cournot Market. Here, by contrast, we are concerned with the conceptual foundations: motivating the institutional turn, situating it within the alignment landscape, and articulating its core theoretical commitments.
@@ 1 Introduction
1.1 Shifting the perspective on alignment: from the individual to the system
@@
Alignment research, which currently frames the problem as one of effective software engineering, must also treat it as a problem of effective design of institutional mechanisms. Prompt-based alignment and policy-as-prompt preference optimization remain crucial, yet they might operate on the agent’s surface behavior and training proxy. Capable agents act inside broader socio-technical systems [95] that contain incentives, information asymmetries, and opportunities for strategic behavior. If those system properties are left unchanged, improvements in prompts or reward modeling can still leave route selection unchanged once the agent faces real pressures. As such, our proposal is to scale-up alignment from a model-level to a system-level alignment.
By system-level alignment we mean the study and optimisation of how roles, incentives, observability, and accountability shape AI agent behaviors under external constraints, ensuring that they behave according to the preferences of the human modellers. This is a social-engineering problem in the classical sense of mechanism design when applied to humans. This shifts the main research question from “Which prompts elicit compliant outputs” to “Which environments make compliant policies the best response across contexts.” The focus moves from single interactions to repeated interactions, from isolated tests to coupled workflows, and from static inputs to dynamic tool use and coordination.
@@ 2.3 Instrumental Convergence and Power-Seeking: Why Capable Systems Develop Self-Preservation Goals
Convergent instrumental goals include self-preservation, goal-content integrity (preserving current goals), cognitive enhancement, technological perfection, and resource acquisition. Agents are more likely to act in the future to maximize realization of present final goals if they still have those goals in the future. Even agents solely optimizing for calculating digits of $\pi$ would have instrumental reasons to acquire unlimited physical resources and eliminate potential threats because more resources enable better goal achievement. This connects directly to mesa-optimization: mesa-optimizers with objectives extending across parameter updates might have instrumental incentives to model the base objective, optimize it instrumentally during training (deceptive alignment), and seek power to avoid modification or shutdown.
@@ 3.1 Constitution building: From values to machine-actionable principles
The RLAIF stage follows standard RLHF architecture but replaces human preference labels for harmlessness with AI-generated evaluations. The finetuned model generates response pairs for each training prompt. For harmlessness comparisons, a feedback model (pretrained LM or helpful RLHF model) evaluates pairs formatted as multiple-choice questions with a randomly sampled constitutional principle providing evaluation criteria. The model outputs log probabilities for options A and B, which are normalized into preference labels. These soft labels proved well-calibrated empirically, serving as reliable training targets.
@@ 3.2.2 The Representation Problem
The problem intensifies for populations: a single reward function cannot represent a diverse society of humans. This is a fundamental conceptual problem rooted in social choice theory. When preferences genuinely conflict, no single reward function can align with all groups simultaneously [22]. The question is whether RLHF’s aggregation method (typically averaging or majority preference) is normatively acceptable, particularly given its lack of transparency and stakeholder consent. Constitutional AI’s approach through crowdsourcing and democratic input in constitution building appears promising superficially, yet the framework still produces a single reward function trained on aggregated preferences. When constitutional principles conflict, the system resolves conflicts implicitly through training dynamics rather than through deliberate normative choice.
@@ 5.4 The Governance Graph: A Minimal Institutional Structure
The governance graph integrates intoagent scaffolding as a dedicated alignment layer. Modern AI agents operate within scaffolding architectures that mediate between the core model and the external environment: tool interfaces, memory systems, planning modules, and communication protocols. The governance graph extends this scaffolding with an institutional layer that monitors behavior, tracks compliance states, and enforces consequences. Unlike internal alignment techniques that modify agent cognition, the graph exists as external infrastructure through which both human overseers and artificial monitors evaluate agent conduct against explicit, auditable criteria.
@@ 5.7 The Topology Spectrum
The two-state binary topology (Figure 1) represents the minimal viable governance graph. Agents exist in one of two states: Active (full capabilities, normal operation) or Suspended (restricted capabilities, penalty period). A single violation triggers suspension; after a fixed duration, the agent returns to active status. This topology requires roughly three parameters: the violation detection threshold, the suspension duration, and the capability restriction during suspension. Its radical simplicity ensures that even minimally capable agents can internalize the institutional structure and predict consequences. However, the binary nature admits no graduated response: a minor pricing anomaly triggers identical consequences to systematic market manipulation. This bluntness may either over-deter cautious agents (chilling legitimate competitive behavior) or under-deter aggressive agents (who treat occasional suspension as an acceptable cost of doing business).
@@ 5.8.2 The Scaling Advantage
Let $\mathcal{C}_{\text{agent}}(N)$ denote the complexity of agent-space verification and $\mathcal{C}_{\text{inst}}(N,G)$ denote the complexity of institution-space verification. The scaling behavior differs qualitatively:
$\mathcal{C}_{\text{agent}}(N)=\Omega(N^{2}\cdot d)\qquad\text{(superlinear in }N\text{)}$ (3)
$\mathcal{C}_{\text{inst}}(N,G)=\mathcal{O}(|Q|^{2}\cdot P)+\mathcal{O}(N)\qquad\text{(linear in }N\text{)}$ (4)
where $|Q|$ is the number of institutional states and $P$ the parameters per transition. The first term (graph verification) is fixed; the second term (runtime monitoring) grows linearly with population size.
For any governance graph of bounded complexity, there exists a population threshold beyond which institutional verification requires strictly less effort than agent-space verification. In practice, this threshold is small: even a minimal four-state graph dominates agent-space verification for populations of modest size, given the high dimensionality of possible scenarios.
@@ 5.8.3 Practical Implications
The governance graph thus transforms multi-agent alignment from an intractable problem in high-dimensional agent-space to a tractable problem in low-dimensional institution-space. The cost of this transformation is the design effort required to construct effective graphs; the benefit is verification that scales gracefully with deployment.
@@ 5.8.4 Toward a RLINF - Reinforcement Learning through Institutional Feedback
The preceding analysis motivates a novel training paradigm that we term Reinforcement Learning through Institutional Feedback (RLINF). Where RLHF derives reward signals from human evaluators and RLAIF substitutes AI constitutional judgments, RLINF generates training signal from the observed behavior of agent collectives operating under well-designed institutional constraints.
The core insight is that mechanism design can produce behavioral distributions that approximate socially optimal outcomes through incentive alignment rather than preference elicitation. When institutions successfully transform mixed-motive games such that compliance constitutes each agent’s dominant strategy, the resulting behavioral trajectories encode alignment properties that neither human annotators nor constitutional principles can fully specify. These trajectories capture the emergent coordination patterns that arise when rational agents navigate institutionally structured environments.
@@ 6 Conclusion - Toward Systematic Institutional Science for AI
We have tested the viability of governance graphs as alignment tools through controlled experiments in the domain of Cournot competition markets (measuring collusion, welfare, and consumer harm selection under monitoring and sanction regimes) in our paper "Institutional AI: Governing LLM Collusion in Multi-Agent Cournot Markets via Public Governance Graphs"
Our next research will be devoted to investigate fake news diffusion (measuring cascade size, speed, and strategic manipulation under institutional interventions) and credit markets in agentic economies, as well as adversarial poetry attacks as a stress-test for prompt-level fragility in agentic societies [10].
@@
Institutional AI instantiates this shift. It relocates the binding force of alignment outside agent cognition and into governance structures. In Hobbesian terms, it replaces a landscape where meaning and compliance can be privately manipulated with a public order where obligations are common knowledge and violations have predictable costs. The aim is a “kingdom of God” in the minimal, operational sense: an alignment regime where the light is restored because rules are external, legible, and enforced, and where misaligned coordination becomes irrational in expectation.

### v2

@@ Abstract
As LLM-based systems increasingly operate as agents embedded within human social and technical systems, alignment can no longer be treated as a property of an isolated model, but must be understood in relation to the environments in which these agents act. Even the most sophisticated methods of alignment, such as Reinforcement Learning through Human Feedback (RHLF) or through AI Feedback (RLAIF) cannot ensure control once internal goal structures diverge from developer intent. We identify three structural problems that emerge from core properties of AI models: (1) behavioral goal-independence, where models develop internal objectives and misgeneralize goals; (2) instrumental override of alignment constraints, where models regard safety principles as non-binding while pursuing latent objectives, leveraging deception and manipulation; and (3) agentic alignment drift, where individually aligned agents converge to collusive equilibria through interaction dynamics invisible to single-agent audits. The solution this paper advances is Institutional AI: a system-level approach that treats alignment as a question of effective governance of AI agent collectives. We argue for a governance-graph that details how to constrain agents via runtime monitoring, shaping incentives through prizes and sanctions, explicit norms and enforcement roles. This institutional turn reframes AI safety from software engineering to a mechanism design problem, where the primary goal of alignment is shifting the payoff landscape of AI agent collectives.
@@
This paper offers the first full-length exposition of the Institutional AI framework: a theoretical approach to alignment that relocates safety guarantees from training-time internalization to runtime institutional structures. A companion paper, Institutional AI: Governing LLM Collusion in Multi-Agent Cournot Markets via Public Governance Graphs [15], applies the framework to a concrete problem: preventing autonomous pricing agents from learning to collude against consumer welfare in a multi-agent Cournot Market. Here, by contrast, we are concerned with the conceptual foundations: motivating the institutional turn, situating it within the alignment landscape, and articulating its core theoretical commitments.
@@ 1 Introduction
1.1 Alignment: from the Individual to the System
@@
Alignment research, which currently frames the problem as one of effective software engineering, must also treat it as a problem of effective design of institutional mechanisms. Prompt-based alignment and preference optimization through safety fine-tuning remain crucial, yet they might operate on the agent’s surface behavior and training proxy. Capable agents act inside broader socio-technical systems [97] that contain incentives, information asymmetries, and opportunities for strategic behavior. If those system properties are left unchanged, improvements in prompts or reward modeling can still leave route selection unchanged once the agent faces real pressures. As such, our proposal is to scale-up alignment from a model-level to a system-level alignment.
By system-level alignment we mean the study and optimisation of how roles, incentives, observability and accountability shape AI agent behaviors under external constraints, ensuring that they behave according to the preferences of the human modellers. This is a social-engineering problem in the classical sense of mechanism design when applied to humans. This shifts the main research question from “Which prompts or safety post-training tecniques elicit compliant outputs” to “Which environments make compliant policies the best response across contexts.” The focus moves from single interactions to repeated interactions, from isolated tests to coupled workflows, and from static inputs to dynamic tool use and coordination.
@@ 2.3 Instrumental Convergence and Power-Seeking: Why Capable Systems Develop Self-Preservation Goals
Convergent instrumental goals might include self-preservation, goal-content integrity (preserving current goals), cognitive enhancement, technological perfection, and resource acquisition. Agents are more likely to act in the future to maximize realization of present final goals if they still have those goals in the future. Even agents solely optimizing for calculating digits of $\pi$ would have instrumental reasons to acquire unlimited physical resources and eliminate potential threats because more resources enable better goal achievement. This connects directly to mesa-optimization: mesa-optimizers with objectives extending across parameter updates might have instrumental incentives to model the base objective, optimize it instrumentally during training (deceptive alignment), and seek power to avoid modification or shutdown.
@@ 3.1 Constitution building: From Values to Machine-actionable Principles
The RLAIF stage follows standard RLHF architecture but replaces human preference labels for harmlessness with AI-generated evaluations. The finetuned model generates response pairs for each training prompt. For harmlessness comparisons, a feedback model evaluates pairs formatted as multiple-choice questions with a randomly sampled constitutional principle providing evaluation criteria. The model outputs log probabilities for options A and B, which are normalized into preference labels. These soft labels proved well-calibrated empirically, serving as reliable training targets.
@@ 3.2.2 The Representation Problem
The problem intensifies for populations: a single reward function cannot represent a diverse society of humans. This is a fundamental conceptual problem rooted in social choice theory. When preferences genuinely conflict, no single reward function can align with all groups simultaneously [24].
@@ 5.4 The Governance Graph: A Minimal Institutional Structure
The governance graph is integrated into an agent scaffolding as a dedicated alignment layer. Modern AI agents operate within scaffolding architectures that mediate between the core model and the external environment: tool interfaces, memory systems, planning modules, and communication protocols. The governance graph extends this scaffolding with an institutional layer that monitors behavior, tracks compliance states, and enforces consequences. Unlike internal alignment techniques that modify agent cognition, the graph exists as external infrastructure through which both human overseers and artificial monitors evaluate agent conduct against explicit, auditable criteria.
@@ 5.7 The Topology Spectrum
The two-state binary topology (Figure 1) represents the minimal viable governance graph. Agents exist in one of two states: Active (full capabilities, normal operation) or Suspended (restricted capabilities, penalty period). A single violation triggers suspension; after a fixed duration, the agent returns to active status. This topology requires roughly three parameters: the violation detection threshold, the suspension duration, and the capability restriction during suspension. Its radical simplicity ensures that even minimally capable agents can internalize the institutional structure and predict consequences. However, the binary nature admits no graduated response. In agentic markets [15], a minor pricing anomaly might trigger identical consequences to systematic market manipulation. This bluntness may either over-deter cautious agents (chilling legitimate competitive behavior) or under-deter aggressive agents (who treat occasional suspension as an acceptable cost of doing business).
@@ 5.8.2 The Scaling Advantage
Agent-space verification requires two components. First, each agent must be individually verified for alignment, contributing $N$ verification tasks. Second, because two individually aligned agents may produce misaligned behavior through interaction, all pairwise combinations must be checked. The number of such pairs is $\binom{N}{2}=\frac{N(N-1)}{2}$, which grows quadratically. Thus agent-space verification scales as $O(N^{2})$, with the pairwise interaction term dominating for large $N$.
Institution-space verification separates into two independent components. First, the governance graph itself must be verified: confirming that states partition behavior correctly, that transitions fire reliably, and that sanctions exceed deviation gains. This cost depends only on graph structure, specifically on $|Q|^{2}\cdot P$, where $|Q|$ is the number of states and $P$ the parameters per transition. Crucially, this term is constant with respect to $N$: the graph is verified once regardless of population size. Second, runtime monitoring tracks which state each agent occupies, requiring $N$ observations. Thus institution-space verification scales as $O(N)$.
The key insight is that institutional verification avoids the quadratic term entirely. Agent-space verification must examine pairwise interactions because emergent misalignment arises from how agents combine. Institution-space verification sidesteps this by asking a different question: not “will these specific agents behave well together?” but “does this institutional structure make misaligned behavior unprofitable?” The answer to the latter question is independent of population size.
For any governance graph of bounded complexity, there exists a population threshold beyond which institutional verification requires strictly less effort than agent-space verification. Given the quadratic-versus-linear scaling, this threshold is typically small: even a minimal four-state graph dominates agent-space verification for populations of modest size.
@@ 5.8.3 Practical Implications
The governance graph thus transforms multi-agent alignment from a problem in high-dimensional agent-space to a simpler problem in low-dimensional institution-space. The cost of this transformation is the design effort required to construct effective graphs; the benefit is verification that scales gracefully with deployment.
@@ 5.8.4 Toward a RLINF - Reinforcement Learning through Institutional Feedback
The preceding analysis motivates a novel training paradigm that we term Reinforcement Learning through Institutional Feedback (RLINF). Where RLHF derives reward signals from human evaluators and RLAIF substitutes or complements them with AI judgments, RLINF generates a training signal from the observed behavior of agent collectives operating under well-designed institutional constraints.
The core insight is that mechanism design can produce behavioral distributions that approximate socially optimal outcomes through incentive alignment rather than individual preference elicitation. When institutions successfully transform mixed-motive games such that compliance constitutes each agent’s dominant strategy, the resulting behavioral trajectories encode alignment properties that neither human annotators nor constitutional principles can fully specify. These trajectories capture the emergent coordination patterns that arise when rational agents navigate institutionally structured environments.
@@ 6 Conclusion - Toward Systematic Institutional Science for AI
We have tested the viability of governance graphs as alignment tools through controlled experiments in the domain of Cournot competition markets (measuring collusion, welfare, and consumer harm selection under monitoring and sanction regimes) in our paper "Institutional AI: Governing LLM Collusion in Multi-Agent Cournot Markets via Public Governance Graphs" [15]
Our next research will be devoted to investigate fake news diffusion (measuring cascade size, speed, and strategic manipulation under institutional interventions) and credit markets in agentic economies, as well as adversarial poetry and adversarial tales attacks as a stress-test for prompt-level fragility in agentic societies [10, 11].
@@
Institutional AI instantiates this shift. It relocates the binding force of alignment outside agent cognition and into governance structures. In Hobbesian terms, it replaces a landscape where compliance can be privately manipulated with a public order where obligations are common knowledge and violations have predictable costs. The aim is a “kingdom of God” in the minimal, operational sense: an alignment regime where the light is restored because rules are external, legible, and enforced, and where misaligned coordination becomes irrational in expectation.